RC Meeting, Saturday, January 29 2022

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 10 AM, remote

  • Zoom remote connection information:
 Join ZoomGov Meeting
 Meeting ID: 160 710 0885
 Passcode: 595071
 One tap mobile
 +16692545252,,1607100885# US (San Jose)
 +16468287666,,1607100885# US (New York)
 Dial by your location
       +1 669 254 5252 US (San Jose)
       +1 646 828 7666 US (New York)
       +1 669 216 1590 US (San Jose)
       +1 551 285 1373 US
       833 568 8864 US Toll-free
 Meeting ID: 160 710 0885
 Find your local number: https://jlab-org.zoomgov.com/u/ab2CekLXgQ

Accelerator Status

  1. Jan 28 Day/Swing -> Jan 29 Owl/Day:
    1. Very Good Beam. Day Beam Summary, Swing Beam Summary, Owl Beam Summary

Hall C Status

Current (is what we need, LOL) RC: Carlos Yero

  • Jan 28 Day / Swing (yesterday):
    • We were able to get most of the Q2=5, (SHMS Left) setting, and started changing kinematics early Jan 29 Owl, where things got a little more "interesting"
  • Jan 29 Owl / Day (today):
    • While changing kinematics to Q2 = 2.45 (SHMS Left), the HMS Drive Controller Faulted at 25.035 deg, just shy of the planned 25.02 deg
    • Due to severe weather condition, we were not able to coordinate escorted access during Owl, and have postpone it until after this RC meeting
    • See Jan 29 Day Shift Run Plan for more details of the run plan for today

Current Run Plan

  • Follow detailed instructions of run plan in: Jan/Feb 2022: 8 GeV Run Plan
  • Current Kinematics: Q2=2.45, W=3.20, x=0.21, low epsilon (SHMS left)
    • SHMS settings: 8.16 deg, +6.265 GeV
    • HMS settings: 25.035 deg, -1.644 GeV
    • data accumulated: ~100 %, currently taking 1 hr-long dummy run, then will switch back to LH2 and take additional statistics until we make Controlled Access to reset HMS Drive Controller

Current Issues

  • Follow-up on Installed Ion Chamber (IC):
    • To follow-up on the status of the newly installed IC in-between SHMS Q1 and beamline, here are some points that resulted from an e-mail exchange between Mark Jones, Ryan (PD) and Shawn Frierson:
  1. The Ion Chamber is hooked up to an FSD Channel (PV name: IICSHMSPk), and we are constantly monitoring it LIVE from CH --> Ion Chamber Monitoring
  2. Limits for the FSD Channel: Bias: 65535 Threshold: 65535 Max: 39321 It’s on the IONHC02 node. They are set very high.
  • Vertical beam position on target has changed a little: Dave Meekins is not concerned about an offset of this size (<0.5 mm).
  • There is a rotation issue that has come up if moving HMS between 19 to 25 degrees. It usually gets stuck around this range, please be wary of this issue. We will likely skip settings in this range if this drive controller error cannot be solved. For singles it will not be a big deal to skip settings for HMS at those angles. We need a longer access to address this issue.
  • The Raster Coil pair Ax and Ay are off
  • Diffuser limit at 46uA (Note: is this still valid?)
  • The SHMS encoder could drift after the physical rotation is done. If that happens we would need an access to reboot the encode/power cycle the rotation rack: https://logbooks.jlab.org/entry/3974431

Special Requests

  1. Look at the EDTM signal on the scope whenever there is extended beam-down period. (Mark Jones) See log-entry EDTM Signal Check
  • This request was in response to an EDTM study by Jacob Murphy and Richard Trotta in which multiple separate blobs of EDTM rawTDC Time signal were observed in some dedicated EDTM Study runs, whenever the runs were pre-scaled. See log-entries: EDTM Study, EDTM Study Follow-up
  • Brad has looked into this, and posted a detailed log-entry with some thoughts/observations --> EDTM Triple Counting Study

Notices in Effect for Shift Crew

  • The Raster Coil pair Ax and Ay are off. Raster needs to be at 4x4. Raster gui has been updated - can now request 2x2.
  • With LHe alarms it is good to check the cameras for plumes. If no plume and it is below the LO alarm and above LOLO then monitor with a plot and halog the problem.
  • With LNitrogen alarms are not that critical. Just need to start monitoring the problem with plot and Halog.
  • HMS NMR is not locking stably - the alarm is turned off. Please check on the NMR readback (while it is locked) intermittently to make sure the HMS dipole field is not drifting,
  • Please keep an eye on the HMS NMR during shift. Due to the NMR locking instability, the alarm has been been disabled, so slow drifts of the NMR signal may be missed.
  • See note on lower right TV screen for bringing up Zoom in CH for remote 3rd shifts
  • Please use the strip-chart on bottom, left Wall TV to stay informed on trends of the SHMS and HMS transfer line vacuum and the magnets insulating vacuum.

Crew Chiefs, We have permission to raise the Hall C diffuser raster limit to allow running >46uA to Hall C. This limit is normally imposed when one of the two rasters is not operative. Since the hall was able to double the running raster current to get the same size raster on the target as having both rasters, Keith is allowing us to raise the limit. To get to the setting, open the Hall C diffuser interlocks screen. Go to the bottom right corner and open the Expert screen. Just below and to the right of the FSD bits, raise the "Raster 1 OK Limit" to slightly above the highest current the hall wants to run at. Do not change the Raster 2 limit. Make a log entry with screen shots. Make an ABIL entry, System MPS and an expected removal date of 3 Jan. 2022. Paul


  • Social distancing requirements back in place. Please maintain a separation of 6 feet during shifts.
  • See updated guidance on JLab COVID-19 isolation and quarantine protocols: https://www.jlab.org/memo/jlab-covid-19-isolation-and-quarantine-requirements
    • Communicate with Occupational Medicine if you have COVID-19 symptoms, have tested positive for COVID-19, or have been exposed to someone who has tested positive for COVID-19: (jnewman@jlab.org or jbanks@jlab.org).
    • In accordance with updated CDC masking guidelines and DOE mandate, cloth masks are no longer permitted or accepted as approved face coverings on-site. Only KF94, KN95, surgical masks and N95s* may be used.
    • N95 masks may be obtained from Industrial Hygiene and worn under voluntary use while on-site as long as social distancing is followed and you have current MED13 (respirator medical approval). To obtain a N95, contact Jennifer Williams from Industrial Hygiene.
    • When your work REQUIRES you to be within 6 feet of others, then a N95 (or better respirator) is required. To use a N95 for close proximity work, a MED13, SAF200 and fit test are also required. Fit tests may be coordinated by reaching out to Jennifer Williams.

Useful Links